OpenStudy (anonymous):

Write 2x + 3y = 6 in slope-intercept form.

OpenStudy (anonymous):

Rearrange the equation for y in terms of x, such as y(x) = bx + c

OpenStudy (anonymous):

what

OpenStudy (anonymous):

Isolate Y

OpenStudy (anonymous):

y=mx+b m = slope b = intercept.

OpenStudy (anonymous):

Y=-2/3x+2